• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Lindgren

    Log entry time 00:04:52 on June13,2000

    Entry number 43619

    keyword=shift summary

    16:00 Experiment down Q3 quenched at 14:41. Hall A access to diagnose
    problem.
    17:00 Hall A under control access. Q3 problem under study.
    18:00 Hall A under control access. Q3 problem under study.
    18:30 Ed Folts and Javier Gomez are trying to decipher the problem while others watch. Decided to replace IC's in control circuit module. Also set up a fan to blow air directly on circuit board
    20:30 Guys out of Hall A. Going to beam permit. Waiting for Q3 to ramp up to 1600 amps. Waiting to read quenching monitor voltage on meter
    21:00 Q3 quenched again at about an average of 108 volts
    21:13 Went into Hall A. Reset Q3. Ramp up to 1600.
    21:15 Tuning beam for Hall A
    21:35 Start Run number 1990 .Monitor quenching voltage. Start at 75 mv
    21:47 End Run 1990 1.3 M events
    21:47 Start Run 1991 monitor voltage 95 mv
    22:10 End run 1991 1.3M
    22:11 Start run 1992 monitor voltage 104 mv
    22:07 Q3 quenched. monitor voltage 107 mv. End Run 1992, 0.58M events
    22:30 Enter Hall A, Reset Q3, Ramping up Q3.
    22:40 Magnet back up. monitor voltage 76 mv. Waiting for beam.
    22:45 Start Run 1994. Monitor voltage 80 mv
    22:51 monitor 85 mv
    23:05 Run 1994 ended. monitor 95, 1.3M events
    23:06 Start run 1995
    23:13 Q3 quenched, monitor 104, ended run 1995. 0.77 events
    23:34 Enter Hall, Reset Q3, Ramp up Q3, Wait.
    23:45 Q3 up, Beam ready, monitor 90 mv,
    23:46 Start run 1996
    23:47 monitor 93 mv
    00:03 monitor 100 mv